OpenStudy (anonymous):

Select the equations that are parallel and perpendicular to y = 2x − 1 and that pass through the point (4, 1). parallel: y = 2x − 7 perpendicular: y = -1/2x + 3 parallel: y = 2x + 2 perpendicular: y = -1/2x+ 4 1/2 parallel: y = -1/2x − 7 perpendicular: y = 2x + 3 parallel: y = −2x − 7 perpendicular: y = -1/2x − 1

OpenStudy (texaschic101):

parallel lines have the same slope, so we know the slope will be 2. y - y1 = m(x - x1)...using slope 2 and points(4,1) y - 1 = 2(x - 4) y - 1 = 2x - 8 y = 2x - 8 + 1 y = 2x - 7 <-- this is your parallel line perpendicular lines have negative reciprocal slopes which means your slope will be -1/2. y - y1 = m(x - x1)...using slope -1/2 and points (4,1) y - 1 = -1/2(x - 4) y - 1 = -1/2x + 2 y = -1/2x + 2 + 1 y = -1/2x + 3 <-- this is your perpendicular line